The area of a rectangular floor is 72m square. If its width is half of its length, find its perimeter.

Answer by ayhanca(1)

The sides are a and b.
Area = a*b = 72.
a=b/2
(b/2)*b=72
b^2=144
b=12
a=12/2 = 6
Perimeter = 2a+2b = 2(6)+2(12) = 12+24 = 36m